### 23 Jan 2024 v1

# Balancing the AI Strength of Roles in Self-Play Training with Regret ${}^{+}$ [^1] https://github.com/vshallc/rm-opponent.git

Abstract
When training artificial intelligence for games encompassing multiple roles, the development of a generalized model capable of controlling any character within the game presents a viable option. This strategy not only conserves computational resources and time during the training phase but also reduces resource requirements during deployment. training such a generalized model often encounters challenges related to uneven capabilities when controlling different roles. A simple method is introduced based on Regret Matching ${}^{+}$, which facilitates a more balanced performance of strength by the model when controlling various roles.
1 Introduction
The Self-Play algorithm [2, 4] has gained significant prominence in reinforcement learning due to its efficacy in enhancing the strength of intelligence agents in adversarial games. In competitive games with different roles (e.g. characters, factions, etc) with unique play styles and abilities to choose, such as the StarCraft with three ”races” or the King of Fighters with dozens of characters, the development of a general model for intelligent agents that can play all roles is more data-efficient and cost saving than creating numerous specialized models. By training different roles on the same model, it allows these roles share their feature and action spaces and portions of the model weights; however, despite being controlled by the same model, the strength exhibited by each role can differ considerably.
This discrepancy can be attributed to two main reasons: 1) the roles may not be designed perfectly balanced; 2) the procedure of strategy learning is related with the complexity of the roles’ abilities. A role which is relatively stronger or has less complicated moves may take in advantage in training convergence. Once a role $A$ has learned an dominated strategy to role $B$, in the following self-play iteration, probability of $B$ can explore a counter strategy becomes low, since the data sampled from self-play for training $B$ are almost weak to $A$.
In order to avoid the incidents of above, a regret matching based data manipulating algorithm is presented in following sections.
2 Training with Self-play in Competitive Games
Let $I={1,2,...}$ denote the set of roles player can choose in a game and $|I|=N$. Given two roles $i,j\in I$, where $i$ is controlled by the model of the $t$th iteration and $j$ the $(t-1)$th, the probability of $i$ beats $j$ is $p(win_{i>j})$. The role $i$ is stronger than $j$ if $p(win_{i>j})>p(win_{j>i})$ and $p(win_{i>j})>0.5$ if there is no draw in the game.
Let $t\in\mathcal{T}=\{0,1,...,T\}$ denote the $t$th iteration in a self-play training and $\theta_{t}$ denote the model parameter after $t$ iterations, where $t=0$ is the initial state and $\theta_{0}$ represents the untrained model. Each iteration contains a number of training steps, $S_{t}={1,2,...}$ and $s_{t}\in S_{t}$ denotes the $s$th training step in iteration $t$. Then the win rate of $i$ against $j$ at step $s_{t}$ is $p(win_{i>j}|\theta_{s_{t}})$, or $p_{s_{t}}(win_{i>j})$ for the sake of simplicity.
When training a generalized model, data samples are generated from all combinations of the roles. Typically, the proportion or weight, $w_{s_{t}}(i,j)$, of the sampled data for each combination is uniformly set to be $\frac{1}{N^{2}}$. The win rate of model at step $s_{t}$, $\theta_{s_{t}}$ against $\theta_{t-1}$ across all role combinations can be expressed as
$p_{s_{t}}(win_{all})=\sum_{i,j\in I}p_{s_{t}}(win_{i>j})w_{s_{t}}(i,j)$
. In practical scenarios, even when the overall win rate converges, such as $p_{s_{t}}(win_{all})\geq 1-\epsilon$ , there may still be a set of $n$ combinations for which $1-p_{s_{t}}(win_{i>j})\gg\epsilon$ if $N^{2}\gg n$. For instance, in a game of $10$ roles, there are $10010^{2}=$ 100 combinations, the overall win rate could still be $0.9$ if $90$ combinations exhibit win rates of $1.0$, while the remaining $10$ never win during the training process. This outcome would indicate a failure in generalization.
3 Manipulating the Data Distribution with Regret ${}^{+}$
In order to train an impeccable general model, the basic idea to our algorithm is to rise the weight of training data where the model’s performance is subpar. Suppose there is a master player on the self-play side, who can only manipulate the weight $w_{s_{t}}(i,j)$ to achieve a higher win rate of its side. Consequently, the weaker the model $\theta_{s_{t}}$ performs on a specific combination, the greater volume the training data that will be utilized.
Regret Matching (RM) and Regret ${}^{+}$ (${}^{+}$) [1, 3] method allow agents to randomly select an action proportional to the positive regrets. By applying ${}^{+}$ to the ”master player”, it enables the player to adaptive the data distribution in response to the strategy employed by its opponent (the training side).
Let $R$ denotes the regret matrix of size $N\times N$, where $R(i,j)$ represents the accumulated regret value of the match between the role $i$ on the training side and the role $j$ on the self-play side. It is important to note that the regret matrix $R(i,j)$ here only indicates the regret of the self-play side (role $j$). $R$ is initialized by filling a value of $\frac{1}{N^{2}}$ to each element.
Given a match between $i$ and $j$, the return to $j$ is defined as:
$r_{s_{t}}(i,j)=\begin{cases}1,j\ \text{wins}\\ 0,\text{otherwise}\\ \end{cases}$
Once the result of the match is comes out, exponential smoothing is applied to the return value and this averaged result serves as an updated win rate:
$p_{s_{t}}(win_{i>j})=\bar{r}_{s_{t}}(i,j)=\bar{r}_{s_{t}-1}(i,j)*\gamma+r_{s_{t}}(i,j)*(1-\gamma)$
where $\gamma\in[0,1]$ is the smoothing factor.
Then the expect utility is defined as the overall win rate, which can be calculated as:
$\mathbb{E}(p_{s_{t}})=p_{s_{t}}(win_{all})=\sum_{i,j\in I}p_{s_{t}}(win_{i>j})w_{s_{t}-1}(i,j)$
The regret is defined as the difference between the smoothed win rate and the expect win rate:
$\Delta p_{s_{t}}(i,j)=\bar{r}_{s_{t}}(i,j)-\mathbb{E}(p_{s_{t}})$
and the regret matrix is updated as:
$R_{s_{t}}(i,j)=\max(R_{s_{t}-1}(i,j)+\Delta p_{s_{t}}(i,j),0)$
Finally, new weights for selecting combinations in next match is updated as:
$w_{s_{t}}(i,j)=\begin{cases}\frac{1}{N^{2}},\text{if}\ \sum_{i,j\in I}R_{s_{t}}(i,j)=0,\\ \frac{R_{s_{t}}(i,j)}{\sum_{i,j\in I}R_{s_{t}}(i,j)}*(1-\eta)+\frac{1}{N^{2}}*\eta\ \text{otherwise}\\ \end{cases}$
where $\eta$ is a weight factor.
4 An Example
For a game with $N=3$ roles, the regret matrix $R_{0}$, weights $w_{0}$ and win rates $p_{0}$ are initialized as follow:
$R_{0}=\begin{bmatrix}0.01&0.01&0.01\\ 0.01&0.01&0.01\\ 0.01&0.01&0.01\\ \end{bmatrix},w_{0}=\begin{bmatrix}\frac{1}{9}&\frac{1}{9}&\frac{1}{9}\\ \frac{1}{9}&\frac{1}{9}&\frac{1}{9}\\ \frac{1}{9}&\frac{1}{9}&\frac{1}{9}\\ \end{bmatrix}\text{and}\ p_{0}=\begin{bmatrix}0.5&0.5&0.5\\ 0.5&0.5&0.5\\ 0.5&0.5&0.5\\ \end{bmatrix}$
When role $j=2$ beats $i=1$, $r_{1}(1,2)=1$. Let $\gamma=0.9$,
$p_{1}(win_{i>j})=\bar{r}_{1}(1,2)=p_{1}(win_{1>2})*0.9+r_{1}(1,2)*0.1=0.55$
and $p1$ is updated as:
$p_{1}=\begin{bmatrix}0.5&0.5&0.5\\ 0.5&0.5&0.55\\ 0.5&0.5&0.5\\ \end{bmatrix}$
Then the updated expect utility is:
$\mathbb{E}(p_{1})=p_{1}(win_{all})=\sum_{i,j\in I}p_{1}(win_{i>j})w_{1}(i,j)=0.5056$
and the regret is updated as:
$\Delta p_{1}(i,j)=\bar{r}_{1}(i,j)-\mathbb{E}(p_{1})=0.55-0.5056=0.0444$
$R_{1}=\begin{bmatrix}0&0&0\\ 0&0&0.0444\\ 0&0&0\\ \end{bmatrix}$
It is important to note that if $R_{1}$ is directly used as weights, then only the combination of $(1,2)$ can be selected in next round. This is the rationale behind the use of a weight factor.
As a result, the weight for next round is ($\eta=0.1$):
$w_{1}=\begin{bmatrix}0.0111&0.0111&0.0111\\ 0.0111&0.0111&0.0400\\ 0.0111&0.0111&0.0111\\ \end{bmatrix}$
5 Evaluation
The evaluation is employed on a fighting game of 13 characters. Two models are trained through 10 iterations for self-play. One is vanilla self-play while the other is self-play with regret matching ${}^{+}$. On each iteration, the model continues training until convergence is achieved. Table 1 shows the win rate among all role combinations.
Role ID 1 2 3 4 5 6 7 8 9 10 11 12 13
1 0.3 0.1 0.08 0.3 0.33 0.07 0.1 0.18 0.02 0.03 0.16 0 0.32
2 1 0.43 0.35 0.8 0.93 0.96 0.05 0.88 0.68 0.27 0.88 0.18 0.37
3 0.73 0.53 0.6 0.36 0.16 0 0.38 0.97 0.5 0.6 0.93 0.2 0.2
4 0.88 0.43 0.52 0.5 0.15 0.13 0.1 0.26 0.2 0.08 0.17 0.45 0.23
5 0.72 0.05 0.9 0.87 0.44 0.49 0.63 0.9 0.11 0.91 0.67 0.56 0.5
6 0.87 0.06 0.88 0.84 0.69 0.6 0.65 0.86 0.18 0.85 0.84 0.09 0.4
7 0.8 0.77 0.67 0.95 0.5 0.5 0.43 0.96 0.34 0.28 1 0.3 0.73
8 0.78 0.06 0 0.5 0.1 0.18 0.05 0.53 0.77 0.98 0.05 0.04 0.4
9 0.96 0.37 0.2 0.76 0.93 0.72 0.71 0.2 0.57 0.75 0.68 0.02 0.75
10 0.92 0.67 0.33 1 0.12 0.04 0.65 0.1 0.34 0.6 0.53 0.03 0.33
11 0.89 0.2 0.09 0.8 0.16 0.2 0 1 0.37 0.55 0.5 0.12 0.77
12 1 0.76 0.8 0.53 0.33 0.82 0.72 0.93 1 0.96 0.8 0.5 0.24
13 0.8 0.67 0.77 0.72 0.6 0.4 0.28 0.72 0.4 0.62 0.3 0.76 0.38
(a) Self-play, no regret matching
Role ID 1 2 3 4 5 6 7 8 9 10 11 12 13
1 0.45 0.88 0.62 0.93 0.74 0.32 0.59 0.76 0.68 0.16 0.8 0.37 0.18
2 0.39 0.53 0.24 0.39 0.8 0.64 0.77 0.74 0.61 0.8 0.66 0.31 0.68
3 0.54 0.66 0.55 0.59 0.47 0.67 0.63 0.46 0.77 0.5 0.66 0.76 0.37
4 0.04 0.71 0.64 0.63 0.7 0.25 0.12 0.49 0.47 0.06 0.09 0.67 0.22
5 0.3 0.32 0.51 0.3 0.53 0.21 0.15 0.43 0.24 0.12 0.29 0.31 0.26
6 0.68 0.29 0.34 0.56 0.71 0.52 0.78 0.81 0.81 0.9 0.53 0.73 0.87
7 0.31 0.23 0.24 0.78 0.88 0.39 0.48 0.5 0.74 0.46 0.39 0.42 0.14
8 0.19 0.35 0.67 0.48 0.5 0.22 0.4 0.42 0.73 0.43 0.69 0.68 0.24
9 0.37 0.41 0.3 0.51 0.62 0.09 0.42 0.56 0.54 0.45 0.63 0.11 0.4
10 0.7 0.29 0.5 0.88 1 0.14 0.5 0.55 0.6 0.33 0.85 0.74 0.16
11 0.31 0.4 0.32 0.8 0.56 0.36 0.61 0.5 0.2 0.05 0.53 0.23 0.3
12 0.72 0.55 0.24 0.33 0.73 0.31 0.54 0.36 0.86 0.43 0.88 0.68 0.15
13 0.91 0.37 0.7 0.69 0.62 0.17 0.87 0.59 0.68 0.81 0.79 0.88 0.47
(b) Self-play with regret matching
Table 1: The win rates on all role combinations for model with/without regret matching ${}^{+}$
The variance of table 0(a) is $0.0964$, while of table 0(b) is $0.0554$.